情報計測工学 2 単位 (選択)

Instrumentation System  カルンガル, ステファン · ギディンシ · 講師/ 知能情報工学科知能工学講座

【授業目的】人間の五感の働きを，機械に人間同様もしくはそれ以上の能力にて代行させようとする技術が情報計測である．これまでに物理量を検出する多くのセンサと計測技術が開発され，実用化されている．本講義では，物理システムの制御に不可欠な情報計測技術の基礎事項を理解させる．

【授業概要】情報計測の一般的な流れは，センサを用いて対象の状態を表す物理量を電気信号に変換して計算機に取り込み，人間や機械が計測対象の状態を容易に把握できるデータに変換するものである．本講義では，この一連の処理に用いられる技術，すなわち目的に応じて物理現象や化学現象を利用して作られた各種センサの原理，センサ出力を計測するためのアナログ回路と計算機に取り込むための入力インターフェイス，さらにディジタルデータを人間や機械が計測対象の情報を容易に把握できるように変換する技術などに関して，具体例を示しながら平易に解説する．

【キーワード】センサ【先行科目】『電気回路及び演習』 (1.0)，『電子回路』 (1.0)，『マイクロプロセッ

サ』 (1.0)

【関連科目】『電子回路』 (0.5) 【履修要件】電気回路及び演習，電子回路，マイクロプロセッサを履修しておく

ことが望ましい．【履修上の注意】なし【到達目標】情報計測技術の基本的な事項と 3 年次に開講される「システム設計

及び実験」で必要とされる知識を習得することを目標とする．【授業計画】

1. 計測の基礎

2. 光センサと磁気センサ

3. 圧力センサと温度センサ

4. 位置センサと超音波センサ

5. 湿度センサとガスセンサ

6. オペアンプとアナログ演算回路

7. AD 変換器

8. DA 変換器

9. アナログフィルタ

10. 入出力インターフェイス

11. ディジタル計測制御システム 12. 信号の変換

13. 電子計測器 (指示計器，波形表示装置)

14. 電子計測器 (波形解析装置，記録装置)

15. 質問 · 総括

16. 定期試験

【成績評価基準】出席 (10%)，れぽーと，演習，小テスト (50%), 期末試験 (40%). 【教科書】田所嘉昭著「電子計測と制御」森北出版【参考書】山崎弘郎著「センサ工学の基礎」昭晃堂【授業コンテンツ】 http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=215979 【対象学生】開講コースと同学科の夜間主コース学生も履修可能【連絡先】

⇒ カルンガル(088-656-7488, karunga@is.tokushima-u.ac.jp) Mail (オフィス

アワー: 8.30-5.30) 【備考】授業を受ける際には， 2 時間の授業時間毎に 2 時間の予習と 2 時間の復習をしたうえで授業を受けることが，授業の理解と単位取得のために必要である

Instrumentation System 2 units (selection)

Stephen Karungaru·Githinji · Associate Professor / Intelligent Systems, Department of Information Science and Intelligent Systems

Target⟩ General instrumentation system aims at the creation of machines that mimic the five human common senses. Many sensors have already been proposed and put to practical use to achieve this. This lecture, therefore, aims at the understanding of the indispensable basic instrumentation systems necessary to control a physical system.

Outline⟩ The general flow of information measurement is the conversion of a physical state of an object using a sensor into an electric signal, feeding the data into a computer, and the conversion of the data into a form that man and the machine can easily understand. This lecture, simply explains the input interface into an analog circuit, the computer system used to measure the data and the outputs of various sensors made using physical or chemical phenomenon, that is, the technology used in processing the digital data while showing by concrete examples the technology that a machine can easily use to convert and understand the information.

�������⟩ Sensors Fundamental Lecture⟩ “Lecture and Exercise in Electric Circuits”(1.0), “Electronic

Circuits”(1.0), “Microprocessors”(1.0) Relational Lecture⟩ “Electronic Circuits”(0.5) Requirement⟩ To take this course, it is desirable to have finished the above

courses Notice⟩ none Goal⟩ Understand basic information instrumentation technology and acquire the

knowledge needed in the course ”System design and experiment” to be taken in the 3rd year.

Schedule⟩

1. Measurement basics

2. Optical and magnetic sensors

3. Pressure and temperature sensors

4. Proximity and ultrasonic sensors

5. Humidity and gas sensors

6. Operational amplifier and analog computing circuit

7. Analog-to-digital converter

8. Digital-to-analog converter 9. Analog filter

10. I/O interface

11. Digital measurement control system

12. Signal conversion

13. Electronic measuring instrument (indicating instrument and wavy display device)

14. Electronic measuring instrument (wavy, analytical device and recording equipment)

15. Question summary

16. Exam

Evaluation Criteria⟩ Result: Attendance (10%) and reports and continuous

assessment tests (50%), final exam (40%). Textbook⟩ Yoshiaki Tadokoro ”Electronic measurement and control” Reference⟩ ”Basics of sensor engineering” by Kouro Yamazaki

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=215979

Student⟩ Able to be taken by night course student of same department

Contact⟩

⇒ (Dr.801, +81-88-656-7488, karunga@is.tokushima-u.ac.jp) Mail (Office Ho

ur: Tuesday s, 15:00 17:00))

Note⟩ To pass this class and to fully understand each lecture, two hours each for preparation and review are necessary.